# Tapor: 3D Hand Pose Reconstruction with Fully Passive Thermal Sensing for Around-device Interactions

### v1

@@ Abstract.
This paper presents the design and implementation of Tapor, a privacy-preserving, non-contact, and fully passive sensing system for accurate and robust 3D hand pose reconstruction for around-device interaction using a single low-cost thermal array sensor. Thermal sensing using inexpensive and miniature thermal arrays emerges with an excellent utility-privacy balance, offering an imaging resolution significantly lower than cameras but far superior to RF signals like radar or WiFi. The design of Tapor, however, is challenging, mainly because the captured temperature maps are low-resolution and textureless. To overcome the challenges, we investigate the thermo-depth and thermo-pose properties and present a novel physics-inspired neural network design that learns effective 3D spatial representations of potential hand poses. We then formulate the 3D pose reconstruction problem as a distinct retrieval task, enabling precise determination of the hand pose corresponding to the input temperature map. To deploy Tapor on IoT devices, we introduce an effective heterogeneous knowledge distillation method that reduces the computation by 377$\times$. We fully implement Tapor and conduct comprehensive experiments in various real-world scenarios. The results demonstrate the remarkable performance of Tapor, which is further illustrated by four case studies of gesture control and finger tracking. We envision Tapor to be a ubiquitous interface for around-device control and have released the dataset, software, firmware, and demo videos at https://github.com/IOT-Tapor/TAPOR.
copyright: nonedoi: isbn: conference: ; 2025journalyear: ;price:
Figure 1. Overview of Tapor. Tapor reconstructs the 3D hand pose from a temperature map generated by a single low-cost thermal array sensor through a physics-inspired neural network model that can run in real-time on IoT.
@@ 1. Introduction
As the Internet of Things (IoT) devices and edge computing continue to proliferate globally, the need for seamless and efficient interaction has been growing rapidly. Around-device gesture interaction (Nandakumar et al., 2016; Lien et al., 2016) has become an integral part of modern interactive systems, providing an increasingly essential interface for applications such as smart homes, autonomous vehicles, virtual and augmented reality environments, human-robot interaction, and wearables/hearables, etc. Around-device gesture interaction, focusing on close-range sensing, enables users to control many devices within the same environment using a consistent set of touchless gestures. This capability is particularly crucial for resource-constrained IoT devices, which often lack traditional input mechanisms such as touchscreens or keyboards, thereby driving the rapid expansion of the gesture control market, which is projected to exceed \$57 billion by 2033 (Insights, 2024).
@@
Table 1. Contactless hand pose reconstruction tech.
Sensing
Modality
Privacy
Protection
Sensing
Mode
Sample
Device
Power Cost
RF Devices Strong Active
IWR1843
Radar
>2 W $\sim$299\$
Acoustic Devices Weak Active
miniDSP
UMA-8-SP
>20 W $\sim$155\$
RGB/IR Cameras Very Weak
Quasi-
passive
Leap
Motion
>2.5 W $\sim$179\$
Thermal Array Medium
Fully
passive
Melexis
MLX90640
<0.08 W $\sim$20\$
@@
Second, we propose a novel method of heterogeneous knowledge distillation to transfer the physics-inspired model into an extremely lightweight network for IoT deployment. Specifically, rather than involving attention blocks, we constrain the model structure to purely convolutional neural networks (CNNs), which are efficiently supported by edge devices. We also perform feature dimension reduction to minimize computational overhead. To transfer knowledge between models of different structures and feature dimensions, we propose heterogeneous knowledge distillation with a learnable adapter to align structures and feature dimensions, significantly reducing the computation and successfully delivering a tiny version of Tapor, named NanoTapor, for real-time on-device deployment for IoT.
@@ 2.1. Thermal Array
A thermal array is a small sensor comprising a miniature array of thermal detectors arranged in a grid pattern. Each detector captures the thermal radiant power $\Phi$ emitted by objects within its field of view (FOV). The sensor converts the captured thermal radiation into an estimated surface temperature for each target area, resulting in a temperature map as illustrated in Fig. 2a. Designed in a compact size, thermal arrays usually have a resolution at least two orders lower than mainstream thermal cameras and are correspondingly significantly cheaper. For example, Tapor utilizes the Melexis MLX90640BAB thermal array sensor, which features a resolution of 32$\times$24 and costs about US\$20, with a reported temperature accuracy of ±$\pm 1.5^{\circ}C$, a FOV of $35^{\circ}\times 55^{\circ}$, and a frame rate up to 64 Hz. Thanks to the low price and compact size, thermal array sensors emerge as an attractive sensing modality, especially for IoT devices.
@@ 2.2. Investigative Experiment
Our preliminary study involves five participants, four males and one female, who performed various hand poses in front of a thermal array sensor and an RGB camera to collect thermal and RGB sample pairs, as shown in Fig. 2a. In total, over 1000 samples were collected from each participant. To examine the effectiveness of existing RGB-based models, we employ Mediapipe (Zhang et al., 2020), a well-known and widely used model for 2D hand pose estimation. The hand pose is represented by 21 key points, as depicted in Fig. 3a. We divide the dataset into training, validation, and testing sets at a ratio of 7:1:2. For a comprehensive evaluation, we run Mediapipe on 1) the original RGB images (256 $\times$ 256), 2) the downsampled images (32 $\times$ 24), and 3) the temperature maps (32 $\times$ 24). As illustrated in Fig. 2b, the results show that, Mediapipe performs considerably worse on the thermal array data than on RGB images, with a performance drop of up to 100%. This suggests that existing RGB-based pose estimation models have difficulty adapting to thermal array data, highlighting the need for a novel model design.
@@ 2.3. Thermo-Pose Modeling
Figure 5. A structure overview of the Tapor and NanoTapor models.
@@ 3. Physics-inspired Design
This section presents our physics-inspired neural network design, followed by its ultra-lightweight variant for IoT.
@@ 3.2.3. Position Retrieval Decoder
To understand whether the model functions as intended, we inspect the model in the following two aspects: (1) We assess the model’s operation to ascertain whether it concentrates on the hand region (i.e., the context). As illustrated in Fig. 9, the attention map of the cross-attention layer clearly demonstrates the model’s focus on the hand region, signifying that the learned model is context-aware. (2) Moreover, to confirm whether the key point features possess distinct semantic meanings (i.e., each feature corresponding to a specific hand key point), we showcase attention maps for various hand parts generated using the relevant key point features, as depicted in Fig. 10. The attention maps distinctly highlight their target hand regions. Additionally, the attentions to different fingers and the root point consistently emphasize the palm’s center, which aligns with our objective to reconstruct an accurate and coherent 3D hand pose.
(a) Input thermal map.
(b) Attention map.
Figure 9. Attention map of the hand. The attention map of the cross-attention layer highlights the model’s correct focus on the hand region, indicating effective context awareness.
@@
Figure 10. Attention maps of root and fingers. The attention maps demonstrate that the key point encoder effectively interprets the semantic meanings of the input temperature map. Note the RGB images are only for visualization.
@@ 4. Implementation
Hardware: We fully implement Tapor using the Melexis MLX90640BAB thermal array sensor, featuring a $32\times 24$ element array with a $35^{\circ}\times 55^{\circ}$ field of view (FOV). The FOV can be expanded to $75^{\circ}\times 110^{\circ}$ with the MLX90640BAA model. The sensor’s refresh rate, configurable between 0.5 Hz and 64 Hz, is set to 10 Hz for our experiments. NanoTapor runs on an ESP32-S3 board with an ESP32S3-WROOM-1 MCU.
@@
Training: Tapor and NanoTapor models are trained using the Adam optimizer with a learning rate of 0.0001. Tapor is trained with a batch size of 48 for 200 epochs, while NanoTapor uses a batch size of 128 for 1000 epochs. All firmware, code, data, models, and demo videos are available here: https://github.com/IOT-Tapor/TAPOR.
@@ 5.1. Experimental Setup
Dataset: To evaluate Tapor, we build a dataset comprising over 60,000 arbitrary hand pose samples (i.e., 32$\times$24 temperature maps). As shown in Fig. 11, the dataset is gathered in five different environments covering different use cases: an office, a bathroom, a bedroom, a kitchen, and a meeting room. Ten participants are involved in the data collection process, including eight males and two females, with approval from our institution’s IRB. For ground truth data, we use the Leap Motion to capture 3D hand poses, while the RealSense D435 depth camera is utilized to record depth and RGB images for our investigative experiments in §2.2. To train the Tapor and NanoTapor models, we use a subset of data from six users in the office for training, while the rest is reserved for testing. Specifically, the dataset is split into 26,557 training samples, 3,319 validation samples, and 34,520 testing samples.
@@
(a) Latency across devices.
(b) Gesture recognition f1 score.
Figure 14. Tapor vs. NanoTapor.
@@ 5.2. Overall Performance
System complexity: NanoTapor requires only 1.93M of RAM and 0.502M of Flash for operation, well within the capacity of most IoT devices, such as the esp32-s3-devkitc-1 board which boasts 3.28M of RAM and 3.34M of Flash. As demonstrated in Tab. 2, activating the MLX90640 thermal array sensor and enabling the $I^{2}C$ interface increases power consumption by 84 mW. Additionally, running the NanoTapor system adds a further 124 mW, resulting in a total power consumption of 767 mW. This is significantly lower than other 3D hand pose reconstruction systems, such as the Leap Motion, which consumes over 2500 mW. It is worth noting that the majority of the power consumption is attributed to the ESP32S3 microcontroller unit (MCU) board. By transitioning to a lower-power MCU, the overall power consumption of the Tapor system could be further reduced.
@@ 5.3. Robustness Evaluation
(a) Different hand coverings.
(b) Tapor performances.
Figure 22. Impact of the hand covering.
Impact of hand temperature: While human body temperatures are relatively stable, our hand surface temperatures vary depending on the environment. We thus evaluate the impact of hand temperature by putting the hand in hot, warm, cold, and iced water before respective experiments. The results, as Fig. 21, show robust performance with slightly increased MPJPE when the hand temperature gets lower.
Impact of ambient temperature: We evaluate performance across various ambient temperatures. As shown in Fig. 21, MPJPE increases significantly beyond 26°C, while MPJRPE remains largely unaffected. This suggests that ambient temperature impacts depth estimation more than relative pose reconstruction, similar to the effect of thick gloves on thermal radiation (§2.3). Despite this, with the sufficient accuracy of relative hand pose reconstruction, Tapor can still support important applications like gesture control.
@@
Impact of ambient interference: Certain everyday objects, such as electronic devices or heated items, may exhibit higher temperatures than their surroundings, potentially introducing interference to Tapor. However, since Tapor is designed for around-device interaction, the user’s hand typically dominates the sensor’s field of view during operation, effectively minimizing the influence of surrounding objects. To systematically evaluate Tapor’s robustness against ambient interference, we conducted experiments involving various common heat-emitting objects. These included a running laptop, an active display, a hot water bottle placed near the device, as well as the presence of the user’s body and light tubes within the sensor’s field of view. As illustrated in Fig. 21, the results demonstrate that Tapor remains robust and maintains reliable performance even in the presence of such environmental interferences.
@@ 7. Related Works
Signal-based: Wireless, acoustic, and light signals are exploited as more privacy-friendly alternatives for hand pose reconstruction, mainly focusing on gesture recognition (Sun et al., 2018; Patra et al., 2018; Pu et al., 2013; Chi et al., 2018; Li et al., 2022b) or hand/finger tracking (Mao et al., 2019; Li et al., 2016; Lien et al., 2016; Nandakumar et al., 2016; Wang et al., 2016; Sun et al., 2015). The most related works are Aili (Li et al., 2017), which reconstructs hand skeletons using light blockage information, and HandFi (Ji et al., 2023), which employs WiFi signals for 3D hand pose reconstruction. These approaches usually suffer from limitations like bulky hardware setup (Li et al., 2017), poor generalization capability (Ji et al., 2023), etc.
Wearable-based: Involving various sensors attached to the wrist, fingers, or integrated into gloves, these methods are used for gesture recognition (Kato and Takemura, 2016; Laput and Harrison, 2019; Xu et al., 2022; Lu et al., 2024) and hand pose estimation (Wu et al., 2020; Zhou et al., 2022). For example, MagX (Chen et al., 2021b) uses passive magnets for hand tracking, and RoFin (Zhang et al., 2023) employs LED nodes in gloves for fine-grained tracking. However, wearable-based solutions create adherence issues and do not fit into massive IoT.
Differently, Tapor offers a privacy-preserving, accurate, and non-intrusive solution for 3D hand pose estimation.

### v2

@@ Abstract.
This paper presents the design and implementation of TAPOR, a privacy-preserving, non-contact, and fully passive sensing system for accurate and robust 3D hand pose reconstruction for around-device interaction using a single low-cost thermal array sensor. Thermal sensing using inexpensive and miniature thermal arrays emerges with an excellent utility-privacy balance, offering an imaging resolution significantly lower than cameras but far superior to RF signals like radar or WiFi. The design of TAPOR, however, is challenging, mainly because the captured temperature maps are low-resolution and textureless. To overcome the challenges, we investigate thermo-depth and thermo-pose properties, proposing a novel physics-inspired neural network that learns effective 3D spatial representations of potential hand poses. We then formulate the 3D pose reconstruction problem as a distinct retrieval task, enabling accurate hand pose determination from the input temperature map. To deploy TAPOR on IoT devices, we introduce an effective heterogeneous knowledge distillation method, reducing computation by 377$\times$. TAPOR is fully implemented and tested in real-world scenarios, showing remarkable performance, supported by four gesture control and finger tracking case studies. We envision TAPOR to be a ubiquitous interface for around-device control and have open-sourced it at https://github.com/aiot-lab/TAPOR.
3D hand pose, Gesture recognition, Thermal array, Thermal sensing
copyright: acmlicensedjournal: IMWUTjournalyear: 2025journalvolume: 9journalnumber: 2article: 63publicationmonth: 6doi: 10.1145/3729499submissionid: 8707ccs: Computer systems organization Embedded and cyber-physical systemsccs: Computer systems organization ccs: Computer systems organization ccs: Networks Hardware Sensor applications and deployments
Figure 1. Overview of TAPOR. TAPOR reconstructs the 3D hand pose from a temperature map generated by a single low-cost thermal array sensor through a physics-inspired neural network model. Leveraging the distilled NanoTapor model, the system operates in real-time on IoT devices with negligible performance degradation, making it suitable for a wide range of around-device interaction applications.
@@ 1. Introduction
As the Internet of Things (IoT) devices and edge computing continue to proliferate globally, the need for seamless and efficient interaction has been growing rapidly. Around-device gesture interaction (Nandakumar et al., 2016; Lien et al., 2016) has become an integral part of modern interactive systems, providing an increasingly essential interface for applications such as smart homes, autonomous vehicles, virtual and augmented reality environments, human-robot interaction, and wearables/hearables, etc.. Around-device gesture interaction, focusing on close-range sensing, enables users to control many devices in the same environment using a consistent set of touchless gestures. This capability is particularly crucial for resource-constrained IoT devices, which often lack traditional input mechanisms such as touchscreens or keyboards, thereby driving the rapid expansion of the gesture control market, which is projected to exceed \$57 billion by 2033 (Insights, 2024).
@@
Table 1. Contactless hand pose reconstruction techniques.
Sensing Modality Privacy Protection Sensing Mode Sample Device Power Cost
RF Devices Strong Active IWR1843 Radar >2 W $\sim$299\$
Acoustic Devices Weak Active miniDSP UMA-8-SP >20 W $\sim$155\$
RGB/IR Cameras Very Weak Quasi-passive Basler IMX334 >2.7 W $\sim$199\$
Thermal Array Medium Fully Passive Melexis MLX90640 <0.08 W $\sim$20\$
@@
Second, we propose a heterogeneous knowledge distillation approach to transfer the physics-inspired model into an extremely lightweight network for IoT deployment. Specifically, rather than involving attention blocks, we constrain the model structure to purely convolutional neural networks (CNNs), which are efficiently supported by edge devices. We also perform feature dimension reduction to minimize computational overhead. To transfer knowledge between models of different structures and feature dimensions, we propose heterogeneous knowledge distillation with a learnable adapter to align structures and feature dimensions, significantly reducing the computation and successfully delivering a tiny version of TAPOR, named NanoTapor, for real-time on-device deployment for IoT.
@@ 2.1. Thermal Array
A thermal array is a small sensor comprising a miniature array of thermal detectors arranged in a grid pattern. Each detector captures the thermal radiant power $\Phi$ emitted by objects within its field of view (FOV). The sensor converts the captured thermal radiation into an estimated surface temperature for each target area, resulting in a temperature map as illustrated in Fig. 2a. Designed in a compact size (approximately 12 mm in height and 9.5 $\times$ 9.5 mm in footprint), thermal arrays usually have a resolution at least two orders lower than mainstream thermal cameras and are correspondingly significantly cheaper. For example, TAPOR utilizes the Melexis MLX90640BAB thermal array sensor, which features a resolution of 32$\times$24 and costs about US\$20, with a reported temperature accuracy of ±$\pm 1.5^{\circ}C$, a FOV of $35^{\circ}\times 55^{\circ}$, and a frame rate up to 64 Hz. Thanks to the low price and compact size, thermal array sensors emerge as an attractive sensing modality, especially for IoT devices.
@@ 2.2. Investigative Experiment
Our preliminary study involves five participants, four males and one female, who performed various hand poses in front of a thermal array sensor and an RGB camera to collect thermal and RGB sample pairs, as shown in Fig. 2a. In total, over 1000 samples were collected from each participant. To examine the effectiveness of existing RGB-based models, we employ Mediapipe (Zhang et al., 2020), a well-known and widely used model for 2D hand pose estimation. The hand pose is represented by 21 key points, as depicted in Fig. 3a. For a comprehensive evaluation, we retrain and test the MediaPipe model under three input conditions: 1) the original RGB images (256 $\times$ 256), 2) the downsampled RGB images (32 $\times$ 24), and 3) the thermal array temperature maps (32 $\times$ 24). All models were trained and evaluated using the same 7:1:2 train/val/test dataset split. As illustrated in Fig. 2b, the results show that, Mediapipe performs considerably worse on the thermal array data than on RGB images, with a performance drop of up to 100%. This suggests that existing RGB-based pose estimation models have difficulty adapting to thermal array data, highlighting the need for a novel model design.
@@ 2.3. Thermo-Pose Modeling
(a) Temperature vs. depth.
(b) Radiation area vs. depth.
@@
Figure 5. A structure overview of the Tapor and NanoTapor models.
@@ 3. Physics-inspired Design
This section presents our physics-inspired neural network design, followed by its ultra-lightweight variant for mobile and IoT devices.
@@ 3.2.3. Position Retrieval Decoder
To understand whether the model functions as intended, we inspect the model in the following two aspects: (1) We assess the model’s operation to ascertain whether it concentrates on the hand region (i.e., the context). As illustrated in Fig. 9a, the attention map of the cross-attention layer clearly demonstrates the model’s focus on the hand region, signifying that the learned model is context-aware. (2) Moreover, to confirm whether the key point features possess distinct semantic meanings (i.e., each feature corresponding to a specific hand key point), we showcase attention maps for various hand parts generated using the relevant key point features, as depicted in Fig. 9b. The attention maps distinctly highlight their target hand regions. Additionally, the attentions to different fingers and the root point consistently emphasize the palm’s center, which aligns with our objective to reconstruct an accurate and coherent 3D hand pose.
(a) Attention map of the hand.
(b) Attention maps of root and fingers.
Figure 9. Attention maps of TAPOR. (a) The attention map of the cross-attention layer highlights the model’s correct focus on the hand region, indicating effective context awareness. (b) The attention maps demonstrate that the key point encoder effectively interprets the semantic meanings of the input temperature map. Note that the RGB images are included solely for visualization purposes.
@@ 4. Implementation
Hardware: We fully implement TAPOR using the Melexis MLX90640BAB thermal array sensor, featuring a $32\times 24$ element array with a $35^{\circ}\times 55^{\circ}$ FOV. The FOV can be expanded to $75^{\circ}\times 110^{\circ}$ with the MLX90640BAA model. The sensor’s refresh rate, configurable between 0.5 Hz and 64 Hz, is set to 10 Hz for our experiments. NanoTapor runs on an ESP32-S3 board with an ESP32S3-WROOM-1 MCU.
@@
Training: Tapor and NanoTapor models are trained using the Adam optimizer with a learning rate of 0.0001. Tapor is trained with a batch size of 48 for 200 epochs, while NanoTapor uses a batch size of 128 for 1000 epochs. All firmware, code, data, models, and demo videos are available here: https://github.com/aiot-lab/TAPOR.
@@ 5.1. Experimental Setup
Dataset: To evaluate TAPOR, we build a dataset comprising over 60,000 arbitrary hand pose samples (i.e., 32$\times$24 temperature maps). As shown in Fig. 10, the dataset is gathered in five different environments covering different use cases: an office, a bathroom, a bedroom, a kitchen, and a meeting room. Ten participants (8 males, 2 females), aged 21–27, with hand widths from 9.4 to 13.1 cm and lengths from 17.3 to 20.5 cm, contributed to the dataset. All procedures were approved by our institution’s IRB. During each trial, participants freely moved their hands above the upward-facing sensor, rotating and translating them within the FoV to cover diverse angles and positions, typically with palms facing downward in a natural interaction posture. This ensures variation in hand orientation and position while remaining within the effective sensing range. For ground truth data, we use the Leap Motion to capture 3D hand poses, while the RealSense D435 depth camera is utilized to record depth and RGB images for our investigative experiments in §2.2. To train the Tapor and NanoTapor models, we use a subset of data from six users in the office for training, while the rest is reserved for testing. Specifically, the dataset is split into 26,557 training samples, 3,319 validation samples, and 34,520 testing samples.
@@ 5.2. Overall Performance
(a) Latency across devices.
(b) Gesture recognition f1 score.
Figure 13. Tapor vs. NanoTapor.
@@
System complexity: NanoTapor requires only 1.93M of RAM and 0.502M of Flash for operation, well within the capacity of most IoT devices, such as the ESP32S3-devkitc-1 board which boasts 3.28M of RAM and 3.34M of Flash. As demonstrated in Tab. 2, activating the MLX90640 thermal array sensor and enabling the $I^{2}C$ interface increases power consumption by 84 mW. Additionally, running the NanoTapor system adds a further 124 mW, resulting in a total power consumption of 767 mW. This is significantly lower than other 3D hand pose reconstruction systems, such as the Leap Motion, which consumes over 1495 mW. It is worth noting that the majority of the power consumption is attributed to the ESP32S3 microcontroller unit (MCU) board. By transitioning to a lower-power MCU, the overall power consumption of the TAPOR system could be further reduced.
@@
(a) Different hand coverings.
(b) TAPOR performances.
Figure 21. Impact of the hand covering.
@@ 5.3. Robustness Evaluation
Impact of hand temperature: While human body temperatures are relatively stable, our hand surface temperatures vary depending on the environment. To evaluate the impact of hand temperature, participants immersed their hands in hot, warm, cold, and iced water for over 20 seconds before each respective experiment. The results, as Fig. 19, show robust performance with slightly increased MPJPE when the hand temperature gets lower.
Impact of ambient temperature: We evaluate performance under various ambient temperatures to study the effect of environmental thermal radiation. To isolate this factor, participants remained in rooms with different ambient temperatures for over five minutes prior to data collection, allowing the thermal background to stabilize. As shown in Fig. 19, MPJPE increases significantly beyond 26°C, while MPJRPE remains largely unaffected. This suggests that ambient temperature impacts depth estimation more than relative pose reconstruction, similar to the effect of thick gloves on thermal radiation. Despite this, with the sufficient accuracy of relative hand pose reconstruction, TAPOR can still support important applications like gesture control.
@@
Impact of ambient interference: Certain everyday objects, such as electronic devices or heated items, may exhibit higher temperatures than their surroundings, potentially introducing interference to TAPOR. However, since TAPOR is designed for around-device interaction, the user’s hand typically dominates the sensor’s FoV during operation, effectively minimizing the influence of surrounding objects. To systematically evaluate TAPOR’s robustness against ambient interference, we conducted experiments involving various common heat-emitting objects. These included a running laptop, an active display, a hot water bottle placed near the device, as well as the presence of the user’s body and light tubes within the sensor’s FoV. As illustrated in Fig. 19, the results demonstrate that TAPOR remains robust and maintains reliable performance even in the presence of such environmental interferences.
@@ 6. Discussions and Future Works
Better accuracy: It remains an open question whether thermal array-based methods can achieve hand pose estimation accuracy comparable to RGB camera-based systems. Exploring approaches such as stereo thermal arrays or multi-sensor fusion may help close this gap and is an important direction for future work.
Hand orientation failure: Cases in which the palm faces upward (i.e., away from the sensor) are rare but represent a potential failure mode due to limited training data for such scenarios. Since our work focuses on around-device interactions, we primarily consider common postures where the palm naturally faces downward toward the sensor, as is typical in interaction scenarios.
@@ 7. Related Works
Signal-based: Wireless, acoustic, and light signals are exploited as more privacy-friendly alternatives for hand pose reconstruction, mainly focusing on gesture recognition (Sun et al., 2018; Patra et al., 2018; Pu et al., 2013; Chi et al., 2018; Li et al., 2022b) or hand/finger tracking (Mao et al., 2019; Li et al., 2016; Lien et al., 2016; Nandakumar et al., 2016; Wang et al., 2016; Sun et al., 2015). Among these, Aili (Li et al., 2017) reconstructs hand skeletons by capturing the blockage of light from a custom LED-panel-and-photodiode-array setup, while HandFi (Ji et al., 2023) leverages WiFi signals to infer 3D hand poses using the existing communication infrastructure. Despite their creativity, these systems face challenges: Aili relies on bulky, non-integrable hardware not suited for compact IoT environments, while HandFi suffers from weak boundary localization and generalization limitations, which may cause unintended activations in real-world settings.
Wearable-based: Involving various sensors attached to the wrist, fingers, or integrated into gloves, these methods are used for gesture recognition (Kato and Takemura, 2016; Laput and Harrison, 2019; Xu et al., 2022; Lu et al., 2024) and hand pose estimation (Wu et al., 2020; Zhou et al., 2022). For instance, MagX (Chen et al., 2021b) enables untethered hand tracking by detecting magnetic fields from passive magnet rings worn on fingers, while RoFin (Zhang et al., 2023) employs custom gloves with embedded LED nodes to achieve fine-grained, real-time 3D pose reconstruction. In a similar vein, Pyro (Gong et al., 2017) introduces a compact, low-power thumb-tip gesture recognition system using pyroelectric sensors. Its design demonstrates robustness across lighting conditions and motion scenarios, and its small form factor makes it attractive for wearable and mobile use.
However, despite their performance, all these wearable-based systems require users to wear specific devices, which can be intrusive or uncomfortable over extended periods. This makes them less suitable for ambient or large-scale IoT deployments, where ease of use, unobtrusiveness, and seamless integration are critical.